# OC4-ReID: Occluded Cloth-Changing Person Re-Identification

### v4

@@ I Introduction
Particularly, OC4-ReID is critical for suspects tracking, as suspects often change clothes and utilize obstacles to hide their bodies during escape, which misleads conventional ReID tracking systems. Compared to CC-ReID, OC4-ReID poses two primary challenges: (i) cloth-unrelated discriminatory information reduces due to occlusion; (ii) blurred and low-quality occlusion regions interfere with effective features extraction.
In this paper, we propose two specialized OC4-ReID datasets and a simple yet effective benchmark for OC4-ReID task. Firstly, We adopt a pre-trained human parser [11] to locate the body components of pedestrians in PRCC [12] and LTCC [13] datasets, then randomly select one component to occlude, constructing the OC4-ReID datasets. Secondly, to mitigate the impact of occlusion and mine discriminative local representations, we propose a Train-Test Micro Granularity Screening (T2MGS) module, in which a quality predictor is maintained to evaluate each micro granularity features. During testing, low-quality features are discarded to suppress occlusions. Thirdly, we also propose a Part-Robust Triplet loss (PRT) to supervise partial features, which aims to combine all the body parts and mine truly discriminative representations. In addition, PRT is robust to clothing features in pedestrians’ partial appearances and emphasises the extraction of clothing-unrelated identity information. The main contributions of this paper are summarized as follows:
@@
II OC4-ReID Dataset
@@
III Method
The overall architecture of our method is illustrated in Fig. 3. We first introduce the Train-Test Micro Granularity Screening (T2MGS) module in Section III-A. Following that, we provide a detailed explanation of the Part-Robust Triplet (PRT) loss in Section III-B.
III-A Train-Test Micro Granularity Screening Module
@@
III-B Part-Robust Triplet Loss
@@
where $\mathcal{F}_{\text{max}}(\cdot,\cdot)$ returns the maximum value, and $M$ is the margin in triplet loss, which is set to 0.3 in this paper. For global features $f_{g}$ and $f_{g}^{w}$, we utilize cross-entropy loss [17] as the identity loss to supervise, which are denoted as $\mathcal{L}_{ID}^{g}$ and $\mathcal{L}_{ID}^{p}$. Additionally, similar to [9], we introduce a clothing-based adversarial loss $\mathcal{L}_{CA}$ that collaborates with the clothing classification loss $\mathcal{L}_{C}$ [17] to learn the clothing-unrelated features. The final total loss is calculated as follows:
@@
IV Experiments
IV-A Datasets and Evaluation Metrics
@@ IV-C Compare with State-of-the-art Methods
IV-D Ablation Study of The Proposed Method
@@
IV-E Sensitivity Analysis of Hyper-Parameter Setting
@@
V Conclusion
@@
VI Acknowledgement
This work was supported by National Nature Science Foundation of China (grant no. U21A6003). (Zhihao Chen and Yiyuan Ge contribute equally to this work) (*Corresponding author: Ziyang Wang.)

### v5

@@ I Introduction
Particularly, OC4-ReID is critical for suspects tracking, as suspects often change clothes and utilize obstacles to hide their bodies during escape, which misleads conventional ReID tracking systems. Compared to CC-ReID, OC4-ReID poses two primary challenges: (i) occlusion leads to even less cloth-unrelated pedestrian discrimination information; (ii) blurred and low-quality occlusion regions further interfere with effective features extraction.
In this paper, we propose two specialized OC4-ReID datasets and a simple yet effective benchmark for OC4-ReID task. Firstly, We adopt a pre-trained human parser [10] to locate the body components of pedestrians in PRCC [11] and LTCC [12] datasets, then randomly select one component to occlude, constructing the OC4-ReID datasets. Secondly, to mitigate the impact of occlusion and mine discriminative local representations, we propose a Train-Test Micro Granularity Screening (T2MGS) module, in which a quality predictor is maintained to evaluate each micro granularity features. During testing, low-quality features are discarded to suppress occlusions. Thirdly, we also propose a Part-Robust Triplet loss (PRT) to supervise partial features, which aims to combine all the body parts and mine truly discriminative representations. In addition, PRT is robust to clothing features in pedestrians’ partial appearances and emphasises the extraction of cloth-unrelated identity information. The main contributions of this paper are summarized as follows:
@@
  • The experimental results in both cloth-changing and occluded cloth-changing scenarios demonstrate the effectiveness of the proposed method.
II Related Works
II-A Occluded Person Re-identification (OCC-ReID)
Person ReID matches the target pedestrian from the image gallery, crossing different scenarios and views [27, 28, 29, 30]. All of the above ReID methods assume that the entire body of the pedestrian is visible, ignoring the more challenging case of occlusion. Mainstream OCC-ReID methods utilize pose estimation [32] and body parsing [31, 33] to localize body parts, aiming to mitigate the effects of occlusion. Wang et al. [32] used labeled poses to learn visible local features as well as topological information during the training and testing phases, achieving high detection accuracy. Song et al. [31] utilized human parsing masks to generate attention maps, enabling the extraction of discriminative and robust features in occluded scenarios. However, these methods rely on maintaining an offline pose or segmentation encoder, which limits model flexibility and incurs significant computational overhead.
II-B Cloth-changing Person Re-identification (CC-ReID)
CC-ReID considers a long-term retrieval situation where the pedestrian’s clothing changes. To tackle this, the researchers employed different approaches including GAN [1], 3D information [34], human contour [3], and adversarial loss [8]. Xu et al. [1] enhanced the model’s ability to learn cloth-unrelated features by reconstructing intra-class examples to minimize feature differences while generating adversarial dressing examples across classes. Chen et al. [34] introduced 3D Shape Learning (3DSL), leveraging 3D body reconstruction as an auxiliary task to extract texture-insensitive 3D shape features for CC-ReID. Cui et al. [3] achieved controlled reconstruction by introducing the cloth-unrelated, cloth-related, and pedestrian contour features. Gu et al. [8] introduced a cloth-based adversarial loss which first trains a cloth classifier to capture cloth-related features, then shifts focus to learning cloth-unrelated features in the second stage.
Although the above methods [1, 3, 8, 34] mitigate clothing variations to a certain extent, they are less effective when additional occlusion disturbances are introduced. Therefore, the aim of this paper is to challenge an even tougher scenario with integrated cloth changes and occlusions.
III OC4-ReID Dataset
@@
IV Method
The overall architecture of our method is illustrated in Fig. 3. We first introduce the Train-Test Micro Granularity Screening (T2MGS) module in Section IV-A. Following that, we provide a detailed explanation of the Part-Robust Triplet (PRT) loss in Section IV-B.
IV-A Train-Test Micro Granularity Screening Module
@@
IV-B Part-Robust Triplet Loss
@@
where $\mathcal{F}_{\text{max}}(\cdot,\cdot)$ returns the maximum value, and $M$ is the margin in triplet loss, which is set to 0.3 in this paper. For global features $f_{g}$ and $f_{g}^{w}$, we utilize cross-entropy loss [16] as the identity loss to supervise, which are denoted as $\mathcal{L}_{ID}^{g}$ and $\mathcal{L}_{ID}^{p}$. Additionally, similar to [8], we introduce a clothing-based adversarial loss $\mathcal{L}_{CA}$ that collaborates with the clothing classification loss $\mathcal{L}_{C}$ [16] to learn the cloth-unrelated features. The final total loss is calculated as follows:
@@
V Experiments
V-A Datasets and Evaluation Metrics
@@ V-C Compare with State-of-the-art Methods
V-D Ablation Study of The Proposed Method
@@
V-E Sensitivity Analysis of Hyper-Parameter Setting
@@
Figure 5: (a) and (b) are the Grad-CAM visualization of the baseline and our method on PRCC datasets. (c) visualizes retrieval results and part-based similarities/distances.
V-F Visualization
We conduct visualization experiments to compare OC4-ReID with the baseline. In Fig.5 (a) and (b), we present the Grad-CAM [26] visualization results, showing that our method focuses more effectively on discriminative pedestrian appearance, whereas the baseline is frequently influenced by identity-independent fragmented information. Additionally, Fig.5 (c) visualizes part-based similarities. The results demonstrate that the model effectively assigns low scores to occluded and cloth-related areas while allocating high scores to cloth-unrelated discriminative body parts, highlighting the model’s ability to handle occlusion and cloth-changing. As demonstrated in Fig. 6, we utilize t-SNE [27] to visualize the feature distribution of random ten identities of baseline and our method. It can be observed that our method encourages more compact intra-class representations while increasing inter-class discrimination, which further validates the effectiveness of our proposed approach.
Figure 6: t-SNE feature visualization of baseline and our method on PRCC dataset. Different colors represent different identities.
VI Conclusion